OpenStudy (anonymous):

Identify the slope and y-intercept of the following equation : 4x-4y=10 and please show the solution .. Thanks . :))

OpenStudy (hoblos):

the easiest way to find the slope and y-intercept is by finding the slope-intercept form of the equation which is of the form y=ax+b so slope=a and y-intercept =(0,b) here 4x-4y=10 --> 4y=4x-10 --> y=x - 5/2 slope=1 y-intercept is (0,-5/2)